# RHanDS: Refining Malformed Hands for Generated Images with Decoupled Structure and Style Guidance

### v1

@@ 3.3. The Second Stage: Structure Guidance
Figure 6. The visual comparison of methods with other methods conditioned on different hand styles. The malformed hand is generated based on the original hand, and the structure guidance is reconstructed from the original hand. For our RHanDS, stage-1 only denotes the model subjected to only the first stage of training, stage-2 only denotes the model trained entirely on the multi-style hand-mesh dataset, and RHanDS denotes the model that has been fully trained through the two-stage training.
4. Multi-Style Dataset
@@
We have introduced a hand-refining model that decouples style and structure, along with a two-stage training strategy. In this section, we will introduce a series of datasets purposefully constructed for this method. As shown in Figure 5, Multi-Style Hand Dataset and Multi-Style Hand-Mesh Dataset correspond to the two stages of training, while Multi-Style Malformed Hand-Mesh Dataset serves as the test set. They encompass various data types and styles of hands. Detailed information about these datasets is provided below.
4.1. Multi-Style Paired Hand Dataset
@@
During the first-stage training, to enable the model to refine hands under style guidance while preventing the leakage of hand structure, we propose a dataset consisting of hand pairs in a variety of styles, with each pair having the same style but different gestures. We find these pairs by detecting the different hands of one person. Specifically, we employ the YOLOv8 (ultralytics, 2022) to detect human hands and use  (Contributors, 2020) to detect human pose, which allows us to filter out hands belonging to the same person. This dataset includes 517,096 hand pairs from natural human images that are various in texture, and 2,423 hand pairs from anime images.
4.2. Multi-Style Hand-Mesh Dataset
@@
During the second-stage training, to prevent the model from overfitting to the generation of hands in a single style, we first utilize the SMPL model (Loper et al., 2023) to create hand meshes and get corresponding depth images. Then, we employ the depth-conditioned ControlNet to generate multi-style hand images with text prompts indicating different style categories. We reconstruct the hand mesh from the generated hand images and calculate the MPJPE metric $MPJPE(K,K^{\prime})=\mathbb{E}||K-K^{\prime}||$ to filter out poorly structured hands, where $K$, $K^{\prime}$ are the 2D projection coordinates of the hand joints, corresponding to the SMPL hand mesh and the reconstructed hand mesh. Ultimately, this dataset includes seven different style categories, with 8,000 hand-mesh pairs for each style.
4.3. Multi-Style Malformed Hand-Mesh Dataset
@@
To evaluate the model’s ability to handle multi-style hands, we extract hand images from the HumanArt (Ju et al., 2023) dataset and use stable diffusion to regenerate the hand regions by SDEdit (Meng et al., 2022). For evaluation, the style reference is the malformed hand itself, while the structure reference is reconstructed from the original hands using (Chen et al., 2022). After filtering out the data with reconstruction failures, we collect 1440 pairs in total, consisting of natural styles and thirteen categories of artificial styles.
Method Style Encoder Structure Encoder MPJPE $\downarrow$ Conf. $\uparrow$ FID $\downarrow$ Style Loss (×$\times 100$$\downarrow$
Malformed Dataset - - 40.29 0.86 23.84 2.77
HandRefiner - - 33.27 0.88 38.52 10.13
RHanDS stage-2 stage-2 26.29 0.89 47.76 7.11
RHanDS stage-1* stage-2 29.48 0.88 33.68 5.35
RHanDS stage-1 stage-2* 26.17 0.91 33.56 5.23
RHanDS stage-1 stage-2 27.23 0.89 32.59 5.04
Table 1. Quantitative comparison of the proposed RHanDS with other methods on Multi-Style Malformed Hand-Mesh Dataset. stage-1* denotes that one hand serves as both a style reference and hand image for the first stage training, and stage-2* denotes that only using the SGD dataset for the second stage training.
@@
Table 2. Quantitative comparison of the proposed RHanDS with other methods on datasets (Lu et al., 2023). The method with † represents calculating metrics on the entire image, while the method without † represents calculating metrics only on the hand region.
5. Experiments
5.1. Implementation Details
In order to utilize the generation capability of existing model, the VAE and U-net of our RHanDS are initialized from Stable Diffusion Inpainting v1.5, and the structure encoder is initialized from pretrained depth ControlNet. For the style encoder, CLIP ViT-H/14 (Ilharco et al., 2021) is adopted as the image encoder, and the linear projection network is randomly initialized.
For the first stage, we use our Multi-Style Paired Hand Dataset to train U-net and style encoder for 15k iterations with a learning rate of 1e-5 and with a total batch size of 256 on 8 NVIDIA-A100 GPUs. The default token length $token$ for the style encoder is set to 8. The hand images $x$ are resized to $512\times 512$, and the style references $r_{style}$ are resized to $224\times 224$ with the processing of random horizontal and vertical flips.
For the second stage, we use the combination of Static Gesture Dataset (AI., [n. d.]) (SGD) and our Multi-Style Hand-Mesh Dataset to train structure encoder for 15k iterations with a learning rate of 2e-5 and with a total batch size of 256 on 8 NVIDIA-A100 GPUs. The weight $w$ for the structure encoder is set to 1 during training. The structure references $r_{struc}$ are resized to $512\times 512$, and the depth values of the hand are normalized into [0.2, 1.0].
Additionally, the mask $mask$ we use to indicate the hand region covers the middle 9/16 of the entire image. In order to learn style guidance from the style encoder rather than the remaining skins in the image, we apply a strategy to randomly expand the region indicated by the mask to force the model to repaint all the skin in the image. Meanwhile, we randomly replace the style embedding with a learnable style embedding with a probability of 0.1 and set the mask to the entire image with a probability of 0.5. We adopt noise offset (Guttenberg, 2023) and min-SNR strategy (Hang et al., 2023) for training.
During inference, we add the maxiummaximum (strength = 1.0) to image $x$ and adopt DDPM (Ho et al., 2020) sampler for denoising. We use structure weight $w$ = 0.6, denoising step $steps$ = 25 as default.
5.2. Evaluation Metrics
@@
5.3. Results and Comparison
We conducted a series of qualitative and quantitative analyses to verify the effectiveness of our method.
Comparison between different methods on HandRefiner Dataset. We evaluate our method on two datasets proposed by HandRefiner. The Text2Image Dataset includes 12K images generated with the text descriptions from HAGRID (Kapitanov et al., 2024), and the Image2Image Dataset includes 2K images sampled from HAGRID. Following the setting in HandRefiner, for each image, we extract the hand region and refine the hand, then paste it back to the original image and calculate metrics on the entire image. Meanwhile, we found that calculating metrics on the hand region extracted using a fixed strategy can eliminate the influence of hand proportion and image size on the evaluation metrics. Therefore, we also report the metrics calculated on the hand region. As shown in Table 2, our RHanDS outperform RandRefiner on both datasets.
Comparison between different methods on Multi-Style Malformed Hand-Mesh Dataset. According to the given settings, we use HandRefiner to refine the hands in our multi-style malformed hand-mesh dataset. As shown in Figure 6, the hands refined by HandRefiner are almost fixed in style, which is different from the original hand style. The quantitative experiments in Table 1 show that our RHanDS outperforms HandRefiner in all metrics in terms of the style and structure of refined hands.
@@ 5.3. Results and Comparison
5.4. Ablation Study
[…]
Visualization of hands generated under the guidance of different style and structure references.
Figure 7. Visualization of hands generated under the guidance of different style and structure references.
@@ 5.4. Ablation Study
Effect of two-stage training. Without the first-stage training, we train the entire model on the multi-style hand-mesh dataset with a single hand as both hand image $x$ and style reference $r_{ref}$. As shown in Table 1, this model can achieve performance comparable to the two-stage training model in terms of MPJPE and Conf., but the FID has increased by 15.17, and the style loss has increased by 2.07. Meanwhile, compared with the the model without structure encoder (FID 33.33, Style Loss 5.50), the FID is increased by 14.43 and the style loss is increased by 1.61. These demonstrate that the first-stage training is beneficial for helping the model perceive hand style from style reference.
@@
Effects of different modules. In RHanDS, we use style encoders and structure encoders to guide the style and structure of the hands. To illustrate the effects of these two modules, we separately shield the structure encoder and the style encoder during inference. Specifically, we set the structure weight $w$ to 0 to shield the structure encoder, and using the learnable style embedding rather than extract embedding from the style reference to shield the style encoder. Qualitative analysis in Figure 8 shows that the structure of the hand tends to be deformed without the structure encoder, and the style of the hand is easily changed without the style encoder.
[…]
The visual comparison of the effects of different modules.
Figure 8. The visual comparison of the effects of different modules.
$w$ MPJPE $\downarrow$ Conf. $\uparrow$ FID $\downarrow$ Style Loss (×$\times 100$$\downarrow$
@@
Table 3. Ablation study on the impact of structure weight $w$.
@@
Figure 9. Failure examples. The first row is the malformed hands, the second row is the hands refined by our RHanDS, and the third row is the structure references.
6. Limitations and Discussion
As shown in Figure 9, we have found that the refined results may be unsatisfactory in some cases. 1) In the first column, for specific styles, such as murals, our model might struggle to accurately capture the style of the hand, leading to alterations in texture details. 2) In the second column, our model may not accurately replicate the gloves onto the refined hands for hands-wearing gloves. 3) In the third column, hands holding objects may be unable to maintain their posture after refining. 4) In the fourth column, the hand mesh for structure guidance needs to be manually constructed for malformed hands that have difficulty reconstructing meshes automatically. Despite these limitations, our RHanDS still provides a paradigm for refining malformed hands with decoupled structure and style guidance.
7. Conclusion
@@
In this paper, we introduce a novel hand refining method RHanDS. It decouples hand refining process into style guidance and structure guidance and employs a two-stage training strategy to fully leverage existing data and improve the structure of repainted hands as well as preserve the style in the original image. Furthermore, three datasets designed specifically for this task are built to facilitate research on multi-style hand refining. Qualitative and quantitative experiments demonstrate our method’s superiority in adapting to images of various styles compared with existing methods.
Appendix A Supplementary Details
In this supplementary material, we show the details of our proposed multi-style datasets, a further quantitative evaluation through a user study, and more hand refining results.
Figure 10. The constructing process of our multi-style hand-mesh dataset for the 2nd training stage.
A.1. Details of Multi-Style Dataset
In this subsection, we introduce details about the constructing process and statistics of our multi-style hand-mesh dataset for the second training stage and multi-style malformed hand-mesh dataset for evaluation, respectively.
A.1.1. Multi-Style Hand-Mesh Dataset
Figure 11. Examples of multi-style hand-mesh training dataset for the 2nd training stage.
Figure 10 shows the constructing process of our multi-style hand-mesh dataset. We combine SMPL-H model (Romero et al., 2017b) and stable diffusion model (Rombach et al., 2022) to synthesize the multi-style hand-mesh dataset. We first use SMPL-H model to generate 3d human models with random shapes, hand poses, and body poses. Specifically, we sample the shape and hand pose parameters from random normal distributions and generate the body pose parameters using vposer (Pavlakos et al., 2019). Then, we crop out the hand region from the full human mesh and use a camera with a field-of-view of 45 degrees to render the hand depth images with and without an arm, respectively. Next, We extract canny images from the hand depth images with an arm and feed them into pre-trained canny and depth ControlNet (Zhang et al., 2023) to guide the structure generation. To synthesize hand images of various styles, we manually customize the text prompt for each style to guide the generation process to match each target style. We finally synthesized 10K images for a totaling 7 styles (natural human, oil painting, watercolor, cartoon, sculpture, digital art, garage kit) and filtered out 20% images for each style using the MPJPE metric introduced in the main material. Beyond this, we further collect, crop, and resize 18962 images from SGD dataset (AI., [n. d.]) as one extra 3d rendering style. Figure 11 shows the training dataset examples of the above 8 styles used in stage 2.
A.1.2. Multi-style Malformed Hand-Mesh Dataset
Figure 12. The constructing process of our multi-style malformed hand-mesh for evaluation.
Figure 12 shows the constructing process of our multi-style malformed hand-mesh dataset. We use the HumanArt dataset (Ju et al., 2023) as a reference to construct our dataset for evaluation. Specifically, we first detect, crop, and resize hands from each original image. Then, we employ MobRecon (Chen et al., 2022) to reconstruct hands from the hand images and manually filter out images of badly reconstructed hands. Finally, we use the stable diffusion inpainting model (AI, 2022a) to repaint the center region of each hand to get the corresponding malformed hand image. The HumanArt dataset has human body images of 20 categories, and after filtering, we got 18 categories, totaling 1440 hand images to construct our dataset. The number of images for each style are: 209 dance, 70 watercolor, 38 stained glass, 86 movie, 46 acrobatics, 306 cosplay, 77 sculpture, 112 drama, 86 digital art, 21 shadow play, 51 mural, 24 sketch, 3 kids drawing, 121 oil painting, 26 ink painting, 52 cartoon, 76 garage kits, 36 relief.
A.2. Quantitative Evaluation with User Study
Input: malformed hand image $I_{1}$, refined image of HandRefiner $I_{2}$, refined image of RHanDS $I_{3}$
Output: perfered image $I$
if style of $I_{2}$ is consistent with $I_{1}$ and style of $I_{3}$ is consistent with $I_{1}$ then
       $I\leftarrow$ one of $\{I_{2},I_{3}\}$ with better hand structure
else
       $I^{\prime}\leftarrow$ one of $\{I_{2},I_{3}\}$ with better style consistency
      $I^{\prime\prime}\leftarrow$ one of $\{I_{2},I_{3}\}$ with worse style consistency
      if $I^{\prime}$ has good hand structure then
             $I\leftarrow I^{\prime}$
      else
             if $I^{\prime\prime}$ has good hand structure then
                   $I\leftarrow I^{\prime\prime}$
            else
                   $I\leftarrow None$
             end if
       end if
end if
Algorithm 1 Human evaluation logic
Figure 13. Visual results of user study evaluation.
To further evaluate our RHanDS with current state-of-the-art methods, we add a user study to conduct the subjective comparison with HandRefiner (Lu et al., 2023). We uniformly sample 173 image pairs (3 for kids drawing and 10 for each other style) from our multi-style malformed hand-mesh dataset and show the image pairs of the malformed hand, the hand refined with HandRefiner, and the hand refined with RHanDS to users side-by-side. We ask users to select the preferred hand image with better structure and style consistency following logic shown in Algorithm 1. This logic considers style consistency first and then compares the structure quality of different methods. If refined images of HandRefiner and RHanDS show both bad style consistency and malformed hand structure, users would reject the two methods and select “None” as the final option.
Figure 13 shows the statistics of most human preferences on the three categories for each image pair. For natural human styles (acrobatics, cosplay, dance, drama, movie), the final selection distribution of HandRefiner, RHanDS, and None is 8, 37, 5. And for artificial human styles (garage kits, relief, sculpture, kids drawing, mural, oil painting, shadow play, sketch, stained glass, cartoon, digital art, ink painting, and watercolor), the final preference distribution is 17, 72, 34. Only for one extreme style of shadow play, both HandRefinre and RHanDS failed to refine malformed hands with appropriate style and structure, and most users select “none” (1, 0, 9). Despite that, users prefer our RHanDS over most other styles. These results validate the conclusion that our method can refine malformed hands with more style consistency and structure quality.
A.3. Further Examples
In this subsection, we show more refining examples of RHanDS to demonstrate the general adaption ability of our method. Figure 14 shows the results.
Figure 14. More refining examples of our RHanDS. In each image pair, the malformed hand, refined hand, and corresponding structure guidance are presented sequentially from left to right.

### v2

@@ 3.1 Preliminaries
Figure 3: The visual comparison of RHanDS with other methods on different styles of malformed hands. The malformed hands are generated based on the original hands, and the structure guidances are reconstructed from the original hands.
3.2 The First Stage: Style Guidance
As shown in Figure 3, the existing solutions are limited in perceiving hand style and exhibit poor generalization ability in images with anime, oil painting, and other styles. To solve the problem, we propose to encode the hand style into U-net, enabling the model to refine hands $x$ based on style reference $r_{style}$. Similar to the approach used in IP-Adapter (Ye et al. 2023a), we use CLIP (Radford et al. 2021) image encoder to extract the global image embedding from $r_{style}$, and then use a linear projection network to transform the image embedding into a sequence of features to obtain the style embedding $c_{style}$. Overall, the style encoder $\mathcal{E}_{style}$ consists of a CLIP image encoder and a linear projection network. The style embedding $c_{style}$ that can be formulated as follows:
$c_{style}=\mathcal{E}_{style}(r_{style})$ (2)
Since the style embedding shares the same dimensions as the original text embedding in U-net, we directly substitute the text embedding with the style embedding. As shown in Figure 4, in this stage, the U-net and the style encoder are trained on multi-style paired hands. We randomly choose a hand from the hand pairs and add noise $\epsilon$ to obtain the noisy latent $z_{t}$, and use the other hand to supply style embedding $c_{style}$ to prevent the model from learning hand structure from style guidance. The loss function is as follows:
$\mathcal{L}=\mathbb{E}_{z_{t},\epsilon,t}[||\epsilon-\epsilon_{\theta}(z_{t},c_{style},mask,t)||^{2}_{2}],$ (3)
To build the multi-style paired hand dataset that meets the specified requirements, we collect character images of varying styles, and crop images of two hands from the same character, ensuring that each pair exhibits identical styles but different gestures. The details of the dataset can be found in Section 4.1.
Figure 4: The first stage. In this stage, U-net and style encoder are trained using Multi-Style Paired Hand Dataset for style guidance.
Figure 5: The second stage. In this stage, the structure encoder is trained using Multi-Style Hand-Mesh Dataset for structure guidance.
3.3 The Second Stage: Structure Guidance
Under style guidance, our framework can repaints hands while preserving their original style. However, as shown in Figure 7, it is hard to correct the hand structure stably without explicit structure guidance due to the inherent complexity of hand structure. To address the problem, we introduce the representation of the correct hand structure by estimating the MANO-based (Romero, Tzionas, and Black 2017a) hand mesh from the malformed hand. For malformed hands that are difficult to reconstruct automatically, manual recontruction can be performed using existing tools (ZhUyU1997 2023). Following Lu et al., we render the hand mesh into a depth image, serving as the structure reference $r_{struc}$, to guide the correction of hand structure.
We use the structure encoder $\mathcal{E}_{struc}$ to encode the structure reference $r_{struc}$ into U-net for structure guidance, where $\mathcal{E}_{struc}$ has the same structure as ControlNet (Zhang, Rao, and Agrawala 2023). Moreover, we feed a learnable embedding $c_{l}$ instead of text embedding into the cross-attention layers of $\mathcal{E}_{struc}$. The learnable embedding shares the same dimension as the text embedding and is integrated into the diffusion model similarly to how the text embedding is. Specially, the input of $\mathcal{E}_{struc}$ is consist of structure reference $r_{struc}$, noisy latent $z_{t}$ and learnable embedding $c_{l}$. The output of $\mathcal{E}_{struc}$ is the structure embedding $c_{struc}$ that can be formulated as follows:
$c_{struc}=\mathcal{E}_{struc}(r_{struc},z_{t},c_{l})$ (4)
We feed the structure embedding $c_{struc}^{i}$ into all the outputs $out^{i}$ of the middle block and decoder blocks of U-net through a $1\times 1$ convolution layer $\mathbb{Z}^{i}$ with structure weight $w$:
$out^{i}=out^{i}+w\cdot\mathbb{Z}^{i}(c_{struc}^{i})$ (5)
As shown in Figure 5, in this stage, we freeze the U-net and the style encoder to maintain the capability of style guidance and train the structure encoder on the multi-style hand-mesh dataset. We add noise $\epsilon$ into the hand image to obtain the noisy latent $z_{t}$ and use the structure encoder to obtain the $c_{struc}$ from the corresponding hand mesh.
$\mathcal{L}=\mathbb{E}_{z_{t},\epsilon,t}[||\epsilon-\epsilon_{\theta}(z_{t},c_{style},c_{struc},mask,t)||^{2}_{2}].$ (6)
However, even if we only fine-tune the structure encoder, the performance of style guidance degrades when training with style-limited datasets datasets (Zimmermann and Brox 2017; Chen et al. 2022). Therefore, based on the depth-controlled text-to-image generation pipeline, we use stable diffusion and ControlNet to create a multi-style hand-mesh dataset by using the depth image rendered from random hand meshes. The details of this dataset can be found in Section 4.2.
4 Multi-Style Dataset
@@
We have introduced a hand-refining model that decouples style and structure guidance, along with a two-stage training strategy. In this section, we will introduce a series of datasets purposefully constructed for this method. They encompass various data types and styles of hands. Detailed information about these datasets is provided below.
4.1 Multi-Style Paired Hand Dataset
@@
During the first-stage training, to enable the model to refine hands under style guidance while preventing the leakage of hand structure, we propose a dataset consisting of hand pairs in various styles, each with the same style but different gestures. We find these pairs by detecting the two hands of the same person. Specifically, we employ the YOLOv8 (ultralytics 2022) to detect human hands and use mmpose (Contributors 2020) to detect human pose, which allows us to filter out hands belonging to the same person. This dataset includes 517,096 hand pairs from natural human images with various textures, and 2,423 hand pairs from anime images.
4.2 Multi-Style Hand-Mesh Dataset
@@
During the second-stage training, to prevent the model from overfitting to the generation of hands in a single style, we first utilize the SMPL model with MANO (SMPL-H) (Loper et al. 2015; Romero, Tzionas, and Black 2017b) to create hand meshes and get corresponding depth images. Specifically, we randomly generated the body pose parameters of the SMPL model and hand pose parameters of the MANO model, then combined them to form a complete human model. We cropped out the arm and hand submesh and rendered them into depth images. Then, we employ the depth-conditioned ControlNet to generate multi-style hand images with text prompts indicating different style categories. We reconstruct the hand mesh from the generated hand images and calculate the metric $\textrm{MPJPE}(K,K^{\prime})=\mathbb{E}||K-K^{\prime}||$ to filter out poorly structured hands, where $K$, $K^{\prime}$ are the 2D projection coordinates of the hand joints, corresponding to the MANO hand mesh and the reconstructed hand mesh. Ultimately, this dataset includes 7 different style categories, with 8,000 hand-mesh pairs for each style.
4.3 Multi-Style Malformed Hand-Mesh Dataset
@@
To evaluate the model’s ability to handle multi-style hands, we extract hand images from the HumanArt (Ju et al. 2023) dataset and use SDEdit (Meng et al. 2022) to regenerate the hand regions as corresponding malformed hand images. For evaluation, the style reference is the malformed hand itself, while the structure reference is reconstructed from the original hands using (Chen et al. 2022). After manually filtering out the data with reconstruction failures, we collect 1440 image pairs in total, consisting of 1 natural style and 13 artificial styles.
5 Experiments
5.1 Implementation Details
In order to utilize the generation capability of existing model, the VAE and U-net are initialized from Stable Diffusion Inpainting v1.5, and the structure encoder is initialized from pretrained depth ControlNet. CLIP ViT-H/14 (Ilharco et al. 2021) is adopted as the style encoder, and the linear projection network is randomly initialized.
For the first stage, we use our Multi-Style Paired Hand Dataset to train U-net and style encoder for 15k iterations with a learning rate of 1e-5 and with a total batch size of 256 on 8 NVIDIA-A100 GPUs. The default token length for the style encoder is set to 8. The hand images $x$ are resized to $512\times 512$, and the style references $r_{style}$ are resized to $224\times 224$ with random horizontal and vertical flipping.
For the second stage, we use the combination of Static Gesture Dataset (SGD) (SynthesisAI 2022) and our Multi-Style Hand-Mesh Dataset to train structure encoder for 15k iterations with a learning rate of 2e-5 and with a total batch size of 256 on 8 NVIDIA-A100 GPUs. The weight $w$ for the structure encoder is set to 1 during training. The structure references $r_{struc}$ are resized to $512\times 512$ and the depth values of the hand are normalized into [0.2, 1.0].
Additionally, the mask $mask$ we use to indicate the hand region covers the middle 9/16 of the entire image. In order to learn style guidance from the style encoder rather than the remaining skins in the image, we apply a strategy to randomly expand the region indicated by the mask to force the model to repaint all the skin in the image. Meanwhile, we randomly replace the style embedding with a learnable style embedding with a probability of 0.1 and set the mask to the entire image with a probability of 0.5. We adopt noise offset (Guttenberg 2023) and min-SNR strategy (Hang et al. 2023) for training.
During inference, we add the maximum noise (strength = 1.0) to image $x$ and adopt DDPM (Ho, Jain, and Abbeel 2020) sampler for denoising. We use structure weight $w$ = 0.6, denoising step $steps$ = 25 as default. The refining process takes approximately 3 seconds and requires 4GB GPU memory.
@@ 5.1 Implementation Details
Table 1: Quantitative comparison of the proposed RHanDS with other methods on datasets (Lu et al. 2024). The method with † represents calculating metrics on the entire image, while the method without † represents calculating metrics only on the hand region.
5.2 Evaluation Metrics
@@
Method Style Encoder Structure Encoder MPJPE $\downarrow$ Conf. $\uparrow$ FID $\downarrow$ Style Loss $\downarrow$
Malformed Dataset - - 40.29 0.86 23.84 2.77
HandRefiner - - 33.27 0.88 38.52 10.13
HandRefiner* - - 29.60 0.89 34.12 5.61
RHanDS stage-2 stage-2 26.29 0.89 47.76 7.11
RHanDS stage-1* stage-2 29.48 0.88 33.68 5.35
RHanDS stage-1 stage-2* 26.17 0.91 33.56 5.23
RHanDS stage-1 stage-2 27.23 0.89 32.59 5.04
Table 2: Quantitative comparison of the proposed RHanDS with other methods on the Multi-Style Malformed Hand-Mesh Dataset. Stage-1* denotes using the hand image itself as a style reference for the first stage training, and stage-2* denotes using only the SGD dataset for the second stage training.
Figure 6: Visualization of hands generated under the guidance of different style and structure references.
5.3 Results and Comparison
Comparison between different methods on HandRefiner Dataset. We evaluate our method on two datasets proposed by HandRefiner. The Text2Image Dataset includes 12K images generated with the text descriptions from HAGRID (Kapitanov et al. 2024), and the Image2Image Dataset includes 2K images sampled from HAGRID. Following the setting in HandRefiner, for each image, we extract the hand region and refine the hand, then paste it back to the original image and calculate metrics on the entire image. Meanwhile, we found that calculating metrics on the hand region extracted using a fixed strategy can eliminate the influence of hand proportion and image size on the evaluation metrics. Therefore, we also report the metrics calculated on the hand region. As shown in Table 1, our RHanDS outperform HandRefiner on both datasets.
@@ 5.3 Results and Comparison
Comparison between different methods on Multi-Style Malformed Hand-Mesh Dataset. We compare with HandRefiner on our multi-style malformed hand-mesh dataset to evaluate the generalization ability to various styles. For fair comparison, we also reimplement and train a variant version of handrefiner on our multi-style dataset, which is denoted as HandRefiner*. As shown in Figure 3, the hands generated by HandRefiner are monotonous in terms of style. After training with our multi-style dataset, HandRefiner* can generate hands of various styles while still struggling to generate hands with the consistent color of the original malformed hands. Through decoupled style and structure guidance, our RHanDS refines hand structure while maintaining consistency with the original malformed hand style. Furthermore, the quantitative experiments in Table 2 show that our RHanDS outperforms HandRefiner and HandRefiner* in all the evaluation metrics in terms of style and structure.
Figure 7: The visual comparison of different modules.
5.4 Ablation Study
@@
Effect of two-stage training. Without the first-stage training, we train the entire model on the multi-style hand-mesh dataset with a single hand as both hand image $x$ and style reference $r_{ref}$. As shown in Table 2, this model can achieve performance comparable to the two-stage training model in terms of MPJPE and Conf., but the FID has increased by 15.17, and the style loss has increased by 2.07. Additional experiment shows that compared with the the model without structure encoder (FID 33.33, Style Loss 5.50), the FID is increased by 14.43 and the style loss is increased by 1.61. These demonstrate that the first-stage training is beneficial for helping the model perceive hand style from style reference.
@@
Effects of different modules. In RHanDS, we use style encoder and structure encoder to guide the style and structure of the hands. To illustrate the effects of these two modules, we separately shield the structure encoder and the style encoder during inference. Specifically, we set the structure weight $w$ to 0 to shield the structure encoder and use the learnable style embedding rather than extracting embedding from the style reference to shield the style encoder. Qualitative analysis in Figure 7 shows that the structure of the hand tends to be malformed without the structure encoder, and the style of the hand is easily changed without the style encoder.
$w$ MPJPE $\downarrow$ Conf. $\uparrow$ FID $\downarrow$ Style Loss   $\downarrow$
@@
Table 3: Ablation study on the impact of structure weight $w$.
@@
6 Conclusion
@@
In this paper, we introduce a novel hand refining method RHanDS. It decouples the hand refining process into style and structure guidance and employs a two-stage training strategy to fully leverage existing data, improve the structure of repainted hands, and preserve the style of the original image. Furthermore, three datasets designed specifically for this task are built to facilitate research on multi-style hand refining. Qualitative and quantitative experiments demonstrate our method’s superiority in adapting to images across various styles compared with existing methods.
Acknowledgements
This work is partially supported by National Natural Science Foundation (Grant No. 62072382), Yango Charitable Foundation, and Alibaba Group through Alibaba Research Intern Program.